# Density control of multi-agent swarms via bio-inspired leader-follower plasticity

### v2

@@ Abstract
Collective behavior; density control; leader-follower.
thanks: This paper was not presented at any IFAC meeting. Corresponding authors Mario di Bernardo and Maurizio Porfiri.
@@ 1 Introduction
Across applications, from swarm robotics [1] to search-and-rescue operations [2], leader-follower strategies [3] are used to design the spatial displacement of large swarms, in which controller agents (leaders) are to induce a desired spatial behavior of another group of agents (followers). While often informed by biological systems like fish [4, 5] and birds [6], engineered systems fail to capture two key aspects of their biological counterparts.
First, leaders and followers maintain fixed hierarchies [7] missing behavioral plasticity, that is, the reversible mechanism behind role changes in biological systems due to internal or environmental stimuli [8]. In fish schools, plasticity arises when followers access novel information about predators or food [9]; likewise, in migrating birds, leaders step down from frontal positions in the flock to recover energy [10]. Second, engineered leaders typically act as fixed open-loop signals, unreceptive to the unfolding collective dynamics [11], whereas biological leaders adapt based on collective needs critical to survival, such as in foraging and predator avoidance [12]. Capturing both these aspects would benefit applications in spatial organization of large swarms, granting flexibility to address multiple tasks at lower control efforts and energy costs [13].
@@
Here, we develop a continuum model for a large population of behaviorally plastic leaders and followers solving a density control problem. We draw insight from the literature on reacting mixtures [29], used to describe blood flows [30] and tumors growth [31]. In particular, we model plasticity as a chemical reaction taking place between two fluids, associated with the continuum description of leaders and followers. We assume that plasticity is not common to all agents, where some of them may not be allowed switching their role. Whether or not an agent is a leader or a follower is not distinguishable by the rest of the group. We derive conditions for the existence and stability of the solution, in terms of key parameters that can be part of engineering design. Our approach provides explicit feedback control laws that guarantee convergence of the coupled PDE system through adaptive leader selection.
@@ 2 Problem Statement
Our control objective is to steer the spatial distribution of a large-scale swarm towards a desired configuration through coordinated leader actions and adaptive role assignment. We adopt a continuum approach consisting of coupled convection-diffusion-reaction equations on the unit circle $\mathcal{S}=[-\pi,\pi]$ [^1] For the sake of brevity, throughout the paper, explicit dependencies on spatial coordinate $x$ and time $t$ are omitted when unambiguous.
@@
where $(\cdot)_{t}$ and $(\cdot)_{x}$ denote partial derivatives. The densities $\rho^{L}$, $\rho^{F}$, $\eta^{F}:\mathcal{S}\times\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ are associated with three agents’ subsets: leaders controlled through the periodic velocity field $u:\mathcal{S}\times\mathbb{R}_{\geq 0}\to\mathbb{R}$ that may become followers; plastic followers that may become leaders; and non-plastic followers that cannot be engaged in roles’ switching. Plasticity between leaders and followers take place through the reaction term $q:\mathcal{S}\times\mathbb{R}_{\geq 0}\to\mathbb{R}$; inter- and intra-population interactions are modeled through a cross-convectional non-local term involving the interaction kernel $f$ (soft-core and odd) and the swarm density $\rho=\rho^{L}+\rho^{F}+\eta^{F}$; agent-level stochasticity is captured by diffusion weighted by the positive coefficient $D$.
@@ 3.2 Design of the reacting term
$\rho^{*}:=\rho^{L}-\rho^{F},$ (16)
and $g$ obeys the mass action law
@@
The control action embedded in $u$ (11) and $q$ (15) assumes full knowledge of the swarm dynamics. This easily accommodates swarm robotics scenarios, in which centralized sensing is in place. Moreover, enabling agents’ communication to estimate densities allows for fully distributed implementations [32, 33].
@@
We provide an illustration of how to use Theorem 5. We consider a population of agents interacting via the periodic Morse interaction kernel (long-range attraction, short-range repulsion)
@@
with $i=L,R$ (see [34] for more details). We set $L_{a}=\pi$, $\alpha=2$, $L_{r}=\pi/6$, and $D=0.05$. We select the desired density for the group to be a von Mises, that is,
@@
where we fix mean $\mu=0$ and concentration coefficient $k=1$. $Z$ is chosen so that $\bar{\rho}$ integrates to 1. Using (20), we establish that the fraction of agents allowed to switch role should be larger than $\hat{p}\approx 0.15$ for the desired collective density to be non-negative and integrating to 1.
@@ 3.3 Stability analysis
where we used integration by parts (twice), and exploited the identity $\left[\left(e^{\eta}\right)^{2}\right]_{x}=2e^{\eta}e^{\eta}_{x}$. We can apply the same procedure in (43) to establish
@@
$\left|-2\mathrm{exp}(-Kt)\int_{\mathcal{S}}e^{\eta}\tilde{v}\,\mathrm{d}x\right|\leq 2\mathrm{exp}(-Kt)\|e^{\eta}\tilde{v}\|_{1}\\ \leq 2\mathrm{exp}(-Kt)\|e^{\eta}\|_{2}\|\tilde{v}\|_{2},$ (49)
@@
where $E^{*}$ and $W$ are Laplace transform of $e^{*}$ and $w$, respectively. Given that $a>0$, the application of the final value theorem yields
@@ 5 Agent-based Model
To bridge the gap between our continuum framework and practical implementation, we develop a discrete agent-based model capturing essential PDE dynamics at the individual level. Following the approach in [20], where density dynamics of swarms with transient leaders are considered, the model consists of coupled stochastic differential equations with Gaussian noise and stochasticity due to roles’ switching
@@
Here, $x_{i}$ are positions of leaders and plastic followers, $y_{i}$ are positions of non-plastic followers, $N^{LF}$ and $M$ are their respective numbers, and $W_{i}$ is a standard Wiener process. Control input $u_{i}(t)=u(x_{i},t)$ is computed via spatial sampling, and $\lambda_{i}\in\{0,1\}$ indicates if agent $i$ is a leader ($\lambda=1$) or follower ($\lambda=0$). $\lambda_{i}$ is updated stochastically with rates $\kappa_{LF}(x_{i},t)$ and $\kappa_{FL}(x_{i},t)$ – see (19).
@@ 7 Conclusions
We presented a bio-inspired leader-follower technique for spatial organization of large swarms via density control. Our strategy incorporates two crucial characteristics of natural systems: setting control objectives for the entire collective and introducing behavioral plasticity to tune the leaders-to-follower mass ratio. We derived conditions for existence and global stability of desired solutions, with numerical findings supporting the effectiveness, robustness, and versatility of our approach.

### v3

# Bio-inspired density control of multi-agent swarms via leader-follower plasticity

@@ Abstract
Density control; Leader-follower systems; Mean-field models; Multi-agent systems; PDEs; Swarm robotics.
@@ 1 Introduction
Across applications, from swarm robotics [14] to search-and-rescue operations [37], leader-follower strategies [18] are used for the spatial displacement of large swarms, in which controller agents (leaders) are to induce a desired spatial behavior of another group of agents (followers). While often informed by biological systems like fish [30, 38] and birds [32], engineered systems fail to capture two key aspects of their biological counterparts.
First, leaders and followers maintain fixed hierarchies [26] missing behavioral plasticity, the reversible mechanism behind role changes in biological systems due to internal or environmental stimuli [33]. In fish schools, plasticity arises when followers access novel information about predators or food [11]; likewise, in migrating birds, leaders step down from frontal positions in the flock to recover energy [35]. Second, engineered leaders typically act as fixed open-loop signals, unreceptive to the unfolding collective dynamics [36], whereas biological leaders adapt based on collective needs critical to survival, such as in foraging and predator avoidance [10]. Capturing these aspects would benefit applications in spatial organization of swarms, granting flexibility to address multiple tasks at lower control efforts and energy costs [19].
@@
Here, we develop a continuum model for a large population of behaviorally plastic leaders and followers solving a density control problem. We draw insight from the literature on reacting mixtures [29], used to describe blood flows [28] and tumors growth [16]. In particular, we model plasticity as a chemical reaction taking place between two fluids, associated with the continuum description of leaders and followers. We assume that plasticity is not common to all agents, where some of them may not be allowed switching their role. Whether or not an agent is a leader or a follower is not distinguishable by the rest of the group. We derive conditions for the existence and stability of the solution, in terms of key parameters that can be part of engineering design. Our approach provides explicit feedback control laws that guarantee convergence through adaptive leader selection.
@@ 2 Problem Statement
Our control objective is to steer the spatial distribution of a large-scale swarm towards a desired configuration through coordinated leader actions and adaptive role assignment. We adopt a continuum approach consisting of coupled convection-reaction-diffusion equations on the unit circle $\mathcal{S}=[-\pi,\pi]$(for the sake of brevity, throughout the paper, explicit dependencies on spatial coordinate $x$ and time $t$ are omitted when unambiguous.)
@@
where $(\cdot)_{t}$ and $(\cdot)_{x}$ denote partial derivatives. The densities $\rho^{L}$, $\rho^{F}$, $\eta^{F}:\mathcal{S}\times\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ are associated with three agents’ subsets: leaders controlled through the periodic velocity field $u:\mathcal{S}\times\mathbb{R}_{\geq 0}\to\mathbb{R}$ that may become followers; plastic followers that may become leaders; and non-plastic followers that cannot be engaged in roles’ switching. Plasticity between leaders and followers take place through the reaction term $q:\mathcal{S}\times\mathbb{R}_{\geq 0}\to\mathbb{R}$; inter- and intra-population interactions are modeled through a cross-convectional non-local term involving the interaction kernel $f$ (soft-core and odd) and the swarm density $\rho=\rho^{L}+\rho^{F}+\eta^{F}$; specifically, $(f*\rho)=\int_{\mathcal{S}}f(\{x,y\})\rho(y,t)\,\mathrm{d}y$, where $\{x,y\}$ is the relative position between $x$ and $y$ wrapped on $\mathcal{S}$ (due to periodicity, it is a circular convolution); agent-level stochasticity is captured by diffusion weighted by the coefficient $D>0$.
@@ 3.2 Design of the reacting term
$\rho^{*}:=\rho^{L}-\rho^{F}$ (16)
is the local imbalance between leaders and followers, and $g$ obeys the mass action law
@@
Our control assumes full knowledge of the swarm dynamics. This accommodates swarm robotics scenarios, in which centralized sensing is in place. Moreover, enabling agents’ communication to estimate densities allows for distributed implementations [13, 9].
@@
We illustrate how to use Theorem 5, by considering a population of agents interacting via the periodic Morse interaction kernel (long-range attraction, short-range repulsion)
@@
with $i=L,R$ (see [8] for more details). We set $L_{a}=\pi$, $\alpha=2$, $L_{r}=\pi/6$, and $D=0.05$. We select a von Mises desired density, that is,
@@
where we fix mean $\mu=0$ and concentration coefficient $k=1$. $Z$ is chosen so that $\bar{\rho}$ integrates to 1. Using (20), we establish that the fraction of agents allowed to switch role should be larger than $\hat{p}\approx 0.15$.
@@ 3.3 Stability analysis
where we used integration by parts (twice), and exploited the identity $\left(\left(e^{\eta}\right)^{2}\right)_{x}=2e^{\eta}e^{\eta}_{x}$. We can apply the same procedure in (43) to establish
@@
$\left|-2\mathrm{exp}(-Kt)\int_{\mathcal{S}}e^{\eta}\tilde{v}\,\mathrm{d}x\right|\leq 2\mathrm{exp}(-Kt)\|e^{\eta}\tilde{v}\|_{1}\\ \leq 2\mathrm{exp}(-Kt)\|e^{\eta}\|_{2}\|\tilde{v}\|_{2}.$ (49)
@@
where $E^{*}$ and $W$ are Laplace transform of $e^{*}$ and $w$, respectively. As $a>0$, the final value theorem yields
@@ 5 Agent-based Model
To bridge the gap between our continuum framework and practical implementation, we develop an agent-based model capturing essential PDE dynamics at the individual level. Following the approach in [3], where density dynamics with transient leaders are considered, the model consists of coupled stochastic differential equations with Gaussian noise and stochasticity due to roles’ switching
@@
Here, $x_{i}$ are positions of leaders and plastic followers, $y_{i}$ are positions of non-plastic followers, $N^{LF}$ and $M$ are their respective numbers, and $W_{i}$ is a standard Wiener process. Control $u_{i}(t)=u(x_{i},t)$ is computed via spatial sampling, and $\lambda_{i}\in\{0,1\}$ indicates if agent $i$ is a leader ($\lambda=1$) or follower ($\lambda=0$). $\lambda_{i}$ is updated stochastically with rates $\kappa_{LF}(x_{i},t)$ and $\kappa_{FL}(x_{i},t)$ – see (19).
@@ 7 Conclusions
We presented a bio-inspired leader-follower technique for spatial organization of large swarms via density control. Our strategy incorporates two crucial characteristics of natural systems: setting control objectives for the entire collective and introducing behavioral plasticity to tune the leaders-to-follower mass ratio. We derived conditions for existence and global stability of desired solutions, with numerical findings supporting the effectiveness, robustness, and versatility of our approach. Numerical simulations highlights that our sufficient conditions for stability are conservative.